OpenStudy (anonymous):

What would the instantaneous rate of f(x) = √x + 3 be?

OpenStudy (mathstudent55):

Is it \(y = \sqrt{x} + 3\) or \(y = \sqrt{x + 3} \) ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

would it be 1/6?

OpenStudy (mathstudent55):

\(f(x) = \sqrt{x} + 3\) You need to take the derivative of y with respect to x. \(f'(x) = \dfrac{d}{dx}(\sqrt{x} + 3)\) \(f'(x) = \dfrac{d}{dx}\sqrt{x} + \dfrac{d}{dx}3\) \(f'(x) = \dfrac{d}{dx}x^{\frac{1}{2}} + 0\) \(f'(x) = \dfrac{1}{2} x^{-\frac{1}{2}} + 0\) \(f'(x) = \dfrac{1}{2x^{\frac{1}{2}}}\) \(f'(x) = \dfrac{1}{2\sqrt{x}}\) Now if you are told a specific x value at which you need the rate of change, plug into \(f'(x) = \dfrac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

I haven't learned derivatives yet gaaaah

OpenStudy (mathstudent55):

Then what have you learned already that allows you to find the instataneous rate of change?

OpenStudy (anonymous):

Average R.O.C = f(x) - f(x)/x And then we assume that x is pretty much 0 at the end because it will be so close to 0.. Just learned it today..

OpenStudy (anonymous):

Should I take a picture of the note?

OpenStudy (mathstudent55):

Are you asked a specific x value to find the r.o.c. for?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and assume x is 0 because it is pretty much 0..

OpenStudy (anonymous):

Does it help if i tell you that we also use it to find slope of tangent?

OpenStudy (mathstudent55):

I just don't know this way of doing it, sorry.

OpenStudy (anonymous):

No problem :) Thanks a bunch :)

OpenStudy (agent0smith):

\[\Large \lim_{h \rightarrow 0} \frac{ f(a+h) - f(a) }{ h }\]is it something like that?